Find the instantaneous rate of change of f(x)=x^2-8/3x+4 at x=-3
Answer by Alan3354(69443)

Find the instantaneous rate of change of f(x)=x^2-8/3x+4 at x=-3
------------
If it's f(x)= (x^2-8)/(3x+4):
f(x) = (x^2-8)*(3x+4)^-1
f'(x) = 2x*(3x+4)^-1 + (x^2-8)*(-1)*(3x+4)^-2 * 3
f'(x) = 2x/(3x+4) - 3(x^2-8)/(3x+4)^2
---
f'(-3) = -6/(-5) - 3*1/25
= 27/25 = 1.08